# Matrix models as tau functions of the UC hierarchy of integrable equations. II

### v1

@@ 2.2.2 Example 2
while $C_{a},\,a>0$ are still arbitrary. The two matrices $C_{a},\,a>0$, the two sets $\mathbf{p}^{(a)}=(p^{(a)}_{1},p^{(a)}_{2},\dots)$ which we denote by $\mathbf{p}$ and the two matrices $X_{a}$ play the role of coupling constants of the model.
@@
This model is associated with the graph that is a segment drawn on the sphere: one edge, two vertices and one face, see Section 3.
@@
2.2.3 Chain matrix models
Without details, we point out that one may use many copies of the integral from Proposition 4, case (i), where for simplicity we put all $\mathbf{p}^{(i)},i=1,\dots,n-1$ be $\mathbf{p}(I_{N})$, which allows one to present the products
$\sum_{\lambda^{(i)},i=0,\dots,n-1}\pi^{(0)}s_{[\lambda^{(i)},\lambda^{(i+1)}]}(X_{i},X_{i+1})\pi^{(n)},$ (76)
and also periodic products as multi-matrix integral associated with spheres glued into a chain. On the other hand, one can show that such matrix integrals can be related to the multi-component UC hierarchy introduced in [57].
@@ 3 Multi-matrix models of complex and unitary matrices associated with embedded graphs
An embedded graph put on an oriented Riemann surface without boundary has only faces that are homeomorphic to a disc. The boundary of each face consists of the sides of ribbon edges. One assigns the positive (counterclockwise) orientation to the boundary of each face. It means that each ribbon edge consists of oppositely directed arrows. Thus, the boundary of the face consists of arrows successively placed next to each other with a positively chosen orientation. Let us number each edge with positive numbers from 1 to $n$, where $n$ is the total number of edges. In the case of sphere, the graph with $F$ faces possesses $V=2+n-F$ vertices. To the sides of edge with number $i$ ($i>0$), we assign numbers $i$ and $-i$, thus, all sides of the graph are numbered with numbers from the set ±$\pm 1,\dots,\pm n$. Let us denote this set as $\mathfrak{N}$, with ${\rm card}\,\mathfrak{N}=2n$. Let us also number the faces of the graph with $1,\dots,F$. By going around each face $f_{i}$ in the positive direction, one obtains a set of numbers from the list above, and such a set is defined up to a cyclic permutation; denote such a set by $\mathfrak{f}_{i}$ and call it face cycle associated with $f_{i}$.
Next, we consider the symmetric group $S_{2n}$, which acts on $\mathfrak{N}$. Consider the element of $S_{2n}$, which is the product of the face cycles:
@@
where, at the left-hand side of this formula, there is a composition of the involution $\sigma^{2}=1$ without fixed points and the product of face cycles. In our notations, $\sigma$ acts as simultaneous transpositions $i\leftrightarrow-i$ for all $i=1,\dots,n$. Evidently, it follows also
@@
This is a combinatorial description of the embedded graph222 In [58], it is written in a slightly different way in terms of numbering the half-edges of the dual graph. drawn on a surface. Formulas (78) and (79) are related to dual graphs.
@@
$d\mu(Z)=e^{-N\mathrm{Tr}ZZ^{\dagger}}\prod_{i,j\leq N}d^{2}Z_{i,j}\,.$ (80)
@@
[…]
@@
Let $\Gamma$ be an embedded graph drawn on the sphere $\mathbb{S}^{2}$ with $n$ band edges and $2n$ sides of these edges, numbered $\mathfrak{N}$ as described in the previous paragraph. Then we consider the equipped graph: to the side of edge numbered $i$ ($i\in\mathfrak{N}$), we assign the matrix $X_{i}$, while to the opposite side of the same edge numbered $-i$, we assign the Hermitian conjugate matrix $X_{-i}:=X^{\dagger}_{i}$. Here, each $X_{i}$ is one of the matrices $Z_{1},\dots,Z_{n_{1}},U_{1},\dots,U_{n_{2}}$. Thus, the matrices $X_{1},\dots,X_{n}$ belong to our mixed ensemble. Such ensembles were introduced in [60]. Some common roots of these constructions can also be found in the long-standing work [59]. The solvable cases for a single matrix (either complex or unitary) were studied earlier in [38].
@@
Consider a cycle $\mathfrak{f}_{i}=(j_{1},j_{2},\dots,j_{\ell_{i}})$, where $\ell_{i}$ is the number of edges of the face $f_{i}$. With each face cycle $(\mathfrak{f}_{i})$, we associate the cycle product
$\mathfrak{f}_{i}\,\to\,{\cal F}_{i}\,:\ \ \ \ \ {\cal F}_{i}(X)=X_{j_{1}}C_{j_{1}}\cdots X_{j_{\ell_{i}}}C_{j_{\ell_{i}}}\,.$ (81)
We call such a product face monodromy because it is the product of matrices obtained by going around the face in the positive direction.
Similarly, consider a cycle $\mathfrak{v}_{i}=(k_{1},\dots,k_{\ell(i)})$, where $\ell_{i}$ is the valency of the vertex $i$. We define vertex monodromy, which is the product of matrices corresponding to the cycle around the vertex:
$\mathfrak{v}_{i}\,\to\,{\cal V}_{i}\,:{\cal V}_{i}(X)=X_{k_{1}}C_{k_{1}}\cdots X_{k_{\ell_{i}}}C_{k_{\ell_{i}}}\,.$ (82)
@@
Having two different orientable surfaces and two embedded graphs, one can mate any selected pair of the Schur functions associated with different graphs using any of $[f\ast g]_{i}$, $i=0,1,2$. To apply $[f\ast g]_{i}$, the matrices should have the block form (20).
@@ Acknowledgements
Our work is partly funded within the state assignment of NRC Kurchatov institute. The work of A. Mironov is also partly supported by the grant of the Foundation for the Advancement of Theoretical Physics and Mathematics “BASIS”. Chuanzhong Li is supported by the National Natural Science Foundation of China under Grant No. 12071237 and Natural Science Research Project of Universities in Anhui Province (grant number 2024AH040202).
Appendix A Measures of matrix ensembles
@@
The Haar measure on $\mathbb{U}_{N}$ in its explicit form is written as
@@
Hermitian matrices.
It is natural [5] to take the following measure on the space $\mathfrak{H}_{N}$ of Hermitian $N\times N$ matrices:
$dM=C_{N}\prod_{i\leq j\leq N}e^{-w\left(\Re M_{i,j}\right)^{2}}d\Re M_{i,j}\prod_{i<j\leq N}e^{-w\left(\Im M_{i,j}\right)^{2}}d\Im M_{i,j}$ (91)
$=c_{N}\prod_{1\leq i<k\leq N}\left(x_{i}-x_{j}\right)^{2}\prod_{i=1}^{N}e^{-wx^{2}_{i}}dx_{i}d_{*}U\,,$
where $w>0$ is a parameter, and $x_{i}$ are eigenvalues of the matrix $M$, while $U$ describes its angular part, $U\in\mathbb{U}_{N}$.
Remark 2.
In many applied problems that use random matrices, it is very convenient to propose that the parameter $w$ is proportional to the size $N$ of the matrices [62, 63, 64].
Appendix B Bilinear equations for UC hierarchy

### v3

@@ 2.2.2 Example 2
[…]
@@
see [11].
This integral is associated with the graph that is a segment drawn on the sphere: one edge, two vertices and one face, see Section 3 and Fig 1c in Appendix D.
@@
Proposition 4 provides sufficient (but not necessary) conditions for the integral (61) to be identified as a UC tau function.
Both examples considered above are related to the mating of two matrix models each of which was actually associated with a graph. In the Example 1 both matrix models can be related to the Fig 1b. In the Example 2 both models which we mate can be realted to the Fig 1c. In the next section, we recall the construction of matrix models related to an arbitrary (embedded) graph presented in [46],[47],[55] and the mating of such models under certain conditions results in the UC tau functions.
@@ 3 Multi-matrix models of complex and unitary matrices associated with embedded graphs
Here we present a general construction of matrix models on graphs with corner matrices (or the same on graphs with source matrices). Corner matrices are free parameters and allows to marry two different models on two different graphs.
From previous studies [55] we know that to obtain a multimatrix model whose partition function can be identified with the tau function KP (and TL), our graph must be plotted on a sphere. Thus, we link two different matrix models corresponding to different graphs on two spheres. And we mate these models, their partition functions and tau functions, obtaining the tau function UC at the output.
@@
An embedded graph put on an oriented Riemann surface without boundary has only faces that are homeomorphic to a disc. The boundary of each face consists of the sides of ribbon edges. One assigns the positive (counterclockwise) orientation to the boundary of each face. It means that each ribbon edge consists of oppositely directed arrows. Thus, the boundary of the face consists of arrows successively placed next to each other with a positively chosen orientation. Let us number each edge with positive numbers from 1 to $n$, where $n$ is the total number of edges. In the case of sphere, the graph with $F$ faces possesses V=2+nFV=2+n-F vertices. To the sides of edge with number $i$ ($i>0$), we assign numbers $i$ and i-i, thus, all sides of the graph are numbered with numbers from the set $\pm 1,\dots,\pm n$. Let us denote this set as $\mathfrak{N}$, with ${\rm card}\,\mathfrak{N}=2n$. Let us also number the faces of the graph with $1,\dots,F$. By going around each face $f_{i}$ in the positive direction, one obtains a set of numbers from the list above, and such a set is defined up to a cyclic permutation; denote such a set by $\mathfrak{f}_{i}$ and call it face cycle associated with $f_{i}$.
Combinatorial description of embedded graphs
can be found, say, in [61], section 1.3.3. We send reader to this wonderful textbook for details. We suggest you take a look at Appendix D for illustrative examples. Let us consider the symmetric group $S_{2n}$, which acts on $\mathfrak{N}$. Consider the element of $S_{2n}$, which is the product of the face cycles:
@@
where, at the left-hand side of this formula, there is a composition of the involution $\sigma^{2}=1$ without fixed points and the product of face cycles. In our notations, $\sigma$ acts as simultaneous transpositions $i\leftrightarrow-i$ for all $i=1,\dots,n$:
$\sigma=(1,-1)(2,-2)\cdots(n,-n)$
which the element of $S_{2n}$ with cycle structure $(2^{n})$.
The simplest example $n=2$ is considered in Appendix D and is related to Fig 1b and Fig 1c.
Remark 2.
To compute the right-hand side of 73 using the transposition action on the product of cycles on the left-hand side, recall the cut-and-join action of transposition on the product of cycles. A couple of simple examples: (i) if we take $n=1$ (one edge) and one cycle ${\mathfrak{f}}=(1,-1)$, we get the “cut” action of transposition: $(1,-1)(1,-1)=(1)(-1)={\mathfrak{v}}_{1}{\mathfrak{v}}_{2}$. The only graph with one edge, one face, and two vertices is the segment, denoted by $\Gamma_{1}$. (ii) Next, let’s take $n=1$, two cycles ${\mathfrak{f}}_{1}=(1)$ and ${\mathfrak{f}}=(-1)$, and get the ”joining” action: $(1,-1)(1)(-1)=(1,-1)={\mathfrak{v}}$. This is the graph $\Gamma_{2}$, which can be represented as the equator on a sphere with a vertex on the equator, $\Gamma_{2}$ is the dual graph of $\Gamma_{1}$.
For completeness, let’s give another example, a little more complicated: we want to describe the graph defined by cycles $a$ and $b$ on the torus. To do this, we consider $n=2$ (two ribbon edges) and the cycle ${\mathfrak{f}}=(1,2,-1,-2)$. Then we first apply transpostion (2,2)(2,-2), breaking $\mathfrak{f}$} into a pair of cycles, and then we apply (1,1)(1,-1) , which will merge these two new cycles into one. We get the cycle $(1,-2,-1,2)={\mathfrak{v}}$, which is a vertex cycle.
Since $\sigma^{2}=1$, we additionaly get
@@
Formulas (73) and (74) are related to dual graphs.
This is a combinatorial description of the embedded graph333 In [61], it is written in a slightly different way in terms of numbering the half-edges of the dual graph. drawn on a surface.
@@
To construct matrix models together with their statistical sum, one can either use the relation (73), forgetting about the graphs (combinatorial approach), or forget about the permutation group and use visualization with graphs (geometric approach). It is a matter of preference. In the combinatorial approach we start from the element in $S_{2n}$ with the cycle structure $(\mathfrak{f}_{1})\cdots(\mathfrak{f}_{F})$ which defines the matrix model, and then thanks to (73) we find $(\mathfrak{v}_{1})\cdots(\mathfrak{v}_{V})$ which defines the partition function.
The way to relate (73) to matrix models is to send $i$ from the set $\{\pm 1,\dots,\pm n\}$ to $C_{i}\in{\rm Mat}_{N\times N}$, and ${\mathfrak{f}}_{a}\to{\cal F}_{a},\,a=1,\dots,F$, ${\mathfrak{v}}_{b}\to{\cal V}_{b},\,b=1,\dots,V$, see (76), (77) and (78) below.
@@
$d\mu(Z)=Ce^{-N\mathrm{Tr}ZZ^{\dagger}}\prod_{i,j\leq N}d^{2}Z_{i,j}\,.\quad\int d\mu(Z)=1$ (75)
@@
In what follows in this section, we consider the mixed ensemble, which contains $n=n_{1}+n_{2}$ matrices $Z_{1},\dots,Z_{n_{1}}$ and $U_{1},\dots,U_{n_{2}}$ with measure $d\Omega_{n_{1},n_{2}}=d\mu(Z_{1},\dots,Z_{n_{1}})d\nu(U_{1},\dots,U_{n_{2}})$. Let us denote the set of these $n_{1}+n_{2}$ matrices by $X$, and the related measure by $dX$. The expectation value of any function $f$ of the entries $\{X_{ij},\,1\leq i,j\leq N\}$ in this set of matrices is defined as
$\langle f\rangle=\int_{\Omega_{n_{1},n_{2}}}f\,dX$
@@
Let $\Gamma$ be an embedded graph drawn on the sphere $\mathbb{S}^{2}$ with $n$ band edges and $2n$ sides of these edges, numbered $\mathfrak{N}$ as described in the previous paragraph. Then we consider the equipped graph: to the side of edge numbered $i$ ($i\in\mathfrak{N}$), we assign the matrix $X_{i}$, while to the opposite side of the same edge numbered i-i, we assign the Hermitian conjugate matrix $X_{-i}:=X^{\dagger}_{i}$. Here, each $X_{i}$ is one of the matrices $Z_{1},\dots,Z_{n_{1}},U_{1},\dots,U_{n_{2}}$. Thus, the matrices $X_{1},\dots,X_{n}$ belong to our mixed ensemble. Such ensembles were introduced in [63]. Some common roots of these constructions can also be found in the long-standing work [62]. The solvable cases for a single matrix (either complex or unitary) were studied earlier in [42].
@@
Consider a cycle $\mathfrak{f}_{i}=(j_{1},j_{2},\dots,j_{\ell_{i}})$, where $\ell_{i}$ is the number of edges of the face $f_{i}$. With each face cycle $(\mathfrak{f}_{i})$, we associate two cycle products. The first one is
$\mathfrak{f}_{i}\,\to\,{\cal F}_{i}\,:\ \ \ \ \ {\cal F}_{i}=C_{j_{1}}\cdots C_{j_{\ell_{i}}}\,.$ (76)
which will call face cycle product (or, the same, face monodromy) and the second is
$\mathfrak{f}_{i}\,\to\,{\cal F}_{i}(X)\,:\ \ \ \ \ {\cal F}_{i}(X)=X_{j_{1}}C_{j_{1}}\cdots X_{j_{\ell_{i}}}C_{j_{\ell_{i}}}\,.$ (77)
which can be called dressed face cycle product, or, the same, dressed face monodromy, because it is the product of matrices obtained by going around the face in the positive direction.
Similarly, consider a cycle $\mathfrak{v}_{i}=(k_{1},\dots,k_{\ell(i)})$, where $\ell_{i}$ is the valency of the vertex $i$. We define vertex cycle product or the same vertex monodromy, which is the product of matrices corresponding to the cycle around the vertex:
$\mathfrak{v}_{i}\,\to\,{\cal V}_{i}\,:{\cal V}_{i}=C_{k_{1}}\cdots C_{k_{\ell_{i}}}\,.$ (78)
and in addition we define dressed vertex cycle product (or, the same, dressed vertex monodromy) as
$\mathfrak{v}_{i}\,\to\,{\cal V}_{i}(X)\,:{\cal V}_{i}(X)=X_{k_{1}}C_{k_{1}}\cdots X_{k_{\ell_{i}}}C_{k_{\ell_{i}}}\,.$ (79)
@@
Having two different orientable surfaces and two embedded graphs, one can mate any selected pair of the Schur functions associated with different graphs using any of $f\ast g]_{i}$, $i=1,2,3$. To apply $f\ast g]_{i}$, the matrices should have the block form (18).
@@
4 Discussion
We propose a way to construct UC-tau functions as the partition function of matrix models. We used integrals over unitary and over complex matrices. In fact, we have a huge family of matrix models with the same partition function. We have limited ourselves to presenting the general construction and simple examples. A more detailed and complete study will be published separately.
@@ Acknowledgements
Our work is partly funded within the state assignment of NRC Kurchatov institute. The work of A. Mironov is also partly supported by the grant of the Foundation for the Advancement of Theoretical Physics and Mathematics “BASIS”. Chuanzhong Li is supported by the National Natural Science Foundation of China under Grant No. 12071237 and Natural Science Research Project of Universities in Anhui Province (grant number 2024AH040202). The work of A.O. is an output of a research project implemented as part of the Basic Research Program at the National Research University Higher School of Economics (HSE University).
Appendix A Partitions. The Schur polynomials [11]
We recall that a nonincreasing set of nonnegative integers $\lambda_{1}\geq\cdots\geq\lambda_{k}\geq 0$, we call partition $\lambda=(\lambda_{1},\dots,\lambda_{l})$, and $\lambda_{i}$ are called parts of $\lambda$. The sum of parts is called the weight $|\lambda|$ of $\lambda$. The number of nonzero parts of $\lambda$ is called the length of $\lambda$, it will be denoted $\ell(\lambda)$. See [11] for details. Partitions will be denoted by Greek letters: $\lambda,\mu,\dots$. The set of all partitions is denoted by $\mathop{\mathrm{P}}\nolimits$.
To define the Schur function $s_{\lambda}$, $\lambda\in\mathop{\mathrm{P}}\nolimits$ at the first step we introduce the set of elementary Schur functions $s_{(m)}$ by
$e^{\sum_{m>0}\frac{1}{m}p_{m}x^{m}}=\sum_{m\geq 0}x^{m}s_{(m)}(\mathbf{p})$ (87)
where the variables $\mathbf{p}=(p_{1},p_{2},p_{3},\dots)$ are called power sum variables. For $\lambda=(\lambda_{1},\lambda_{2},\dots)\in\mathop{\mathrm{P}}\nolimits$ we define
$s_{\lambda}(\mathbf{p})=\mathrm{det}\left[s_{(\lambda_{i}-i+j)}(\mathbf{p})\right]_{i,j>0}$ (88)
Let us introduce notations $\mathbf{p}_{\infty}:=(1,0,0,\dots)$ and $\mathbf{p}(a):=(a,a,a,\dots)$. We have (see Ch. I in [11])
$\frac{s_{\lambda}(\mathbf{p}(a))}{s_{\lambda}(\mathbf{p}_{\infty})}=(a)_{\lambda}:=\prod_{(i,j)\in\lambda}(a-j+i)$ (89)
The product ranges all nodes $(i,j)$ of the Young diagram $\lambda$ [11]. The product on the right-hand side is called the content product, since the number jij-i is called the content of the node with coordinates $(i,j)$.
The Schur functions play the role of the characters of the irreducable reperesentation labeled by $\lambda$ of the linear group.
The following relation is known as the character map relation
$s_{\lambda}(X)=s_{\lambda}(\mathbf{p}_{\infty})\sum_{\mu\atop|\mu|=|\lambda|}\varphi_{\lambda}(\mu)\prod_{i=1}^{\ell(\mu)}\mathrm{tr}\left(X^{\mu_{i}}\right),\quad\mathbf{p}_{\infty}:=(1,0,0,\dots),$ (90)
where $\varphi_{\lambda}(\mu)$ are rational numbers which are expressed in terms of the characters of the symmetric group $S_{d},\,d=|\lambda|$, see [11]. It follows from (90) that $s_{\lambda}(X)$ is a polynomial of degree $d=|\mu|=|\lambda|$ in entries of $X$. Therefore, $s_{\lambda}(X)$ is a polynomial of each entry of matrix $X$ and one can integrate it over entries, as it is written in Section B.
Appendix B Measures of matrix ensembles
@@
The Haar measure on $\mathbb{U}_{N}$ is $d_{*}U=\prod_{i,j}\left(dUU^{-1}\right)_{ij}$ and is also written as
@@
Ginibre ensemble of complex matrices
$d\mu(Z)=Ce^{-N\mathrm{Tr}ZZ^{\dagger}}\prod_{i,j\leq N}d^{2}Z_{i,j}\,.$ (92)
where $C$ is chosen in a way $\int d\mu(Z)=1$. The measure of ensemble of many complex matrices $Z_{1},\dots,Z_{n}$ is the product $\prod_{i}^{n}d\mu(Z_{i})$.
Appendix C Bilinear equations for UC hierarchy
@@
Appendix D Examples of graphs for matrix model
For illustration purposes we use figures from [55].
{arrangedFigure}
[…]
Figure 1a shows a fragment of a graph, which shows three edges with side numbers (i,i),(j,j)(i,-i),(j,-j) and (k,k)(k,-k), as well as two vertices highlighted in yellow. Each side is assigned a random matrix with the same number: $(Z_{i},Z_{-i})$ – to the sides with edge number $|i|$, $(Z_{j},Z_{-j})$ – to the sides of the edge numbered $|j|$, and $(Z_{k},Z_{-k})$ – to the sides with edge number $|k|$. It is assumed that $Z_{-a}=Z_{a}^{\dagger}$, where $a=i,j,k$. If you go around a face in the positive direction, then each corner of the face is assigned the number of the previous side of the edge. In the corners, corner matrices (the same: source matrices of the model) are placed, with the same number as the corner number. In our case, these are matrices $C_{-i}$, following the side i-i, then matrix $C_{k}$, following the side $k$, then matrix $C_{j}$, following the side $j$. We also have the corner matrix $C_{-k}$ following the side k-k, and the corner matrix $C_{-j}$ following the side j-j. The vertex cycles are $\mathfrak{v}_{1}=(k,-j)$ and $\mathfrak{v}_{2}=(j,-i,-k)$, and the corresponding vertex monodromies are $\mathfrak{V}_{1}=C_{k},C_{-j}$ and $\mathfrak{V}_{2}=C_{j},C_{-i},C_{-k}$ respectively. It is not possible to write relation (80) because it applies to the whole graph, not to a fragment. It is possible to write the cycle of only one face: $\mathfrak{f}=(-j,-k)$, since the other faces are not represented. The monodromy of this face is $\mathfrak{F}=C_{-j}C_{-k}$, and the corresponding dressed monodromy (dressed by random matrices) of the face is obtained from $C_{a}\to Z_{a}C_{a}$ and is equal to $\mathfrak{F}(Z)=Z_{-j}C_{-j}Z_{-k}C_{-k}$
The graph in Fig. 1b gives us $\mathfrak{v}=(-1,1)$, and the corresponding vertex monodromy is $\mathfrak{V}=C_{1}C_{-1}$. There are two face cycles $(\mathfrak{f}_{1})=(1)$ and $(\mathfrak{f}_{2})=(-1)$. The corresponding face monodromies are $\mathfrak{F}_{1}=C_{1}$ and $\mathfrak{F}_{2}=C_{-1}$, and the dressed monodromies are $\mathfrak{F}_{1}=Z_{1}C_{1}$ and $\mathfrak{F}_{2}=Z_{-1}C_{-1}$. The involution $\sigma\in S_{2}$ is the cycle (1,1)(1,-1). We have an obvious relation (73): $\sigma\circ(\mathfrak{f}_{1})(\mathfrak{f}_{2})=(-1,1)$, which desribes the ”join action” which unifies the cycles $(\mathfrak{f}_{1})$ and $(\mathfrak{f}_{2})$. In this case, relation (80) takes the form (46).
Let us note that the graph on Fig 1c is dual to the graph on Fig 1b. The relation (81) reads as (9). It is related to the “cut action”: $(-1,1)\circ(-1,1)=(1)(-1)$.
[…]
The graph in Fig. 2a is depicted on a torus. Its only face is the fundamental domain depicted in Fig. 2b. The face cycle is $\mathfrak{f}=(1,2,-1,-2)$, the face monodromy is $(\mathfrak{F})=C_{1}C_{2}C_{-1}C_{-2}$, the dressed face monodromy is $\mathfrak{F}(Z)=Z_{1}C_{1}Z_{2}C_{2}Z_{-1}C_{-1}Z_{-2}C_{-2}$. The vertex cycle is $\mathfrak{v}=(1,-2,-1,2)$. Relation (80) in the case of $Z\in GL_{N}$ has the form
$\int s_{\lambda}\left(\mathfrak{F}(Z)\right)d\mu(Z)=N^{-2|\lambda|}s_{\lambda}(\mathbf{p}_{\infty})^{-2}s_{\lambda}(\mathfrak{V}),\quad\mathfrak{V}=C_{1}C_{-2}C_{-1}C_{2}$
The matrix model associated with this graph is
$\int e^{\sum_{m>0}N\frac{p_{m}}{m}\mathrm{Tr}\left({\mathfrak{F}(Z)}\right)^{m}}d\mu(Z)=\sum_{\lambda\atop\ell(\lambda)\leq N}N^{-2|\lambda|}\frac{s_{\lambda}(N\mathbf{p})s_{\lambda}(C_{1}C_{-2}C_{-1}C_{2})}{s_{\lambda}(\mathbf{p}_{\infty})^{2}}$ (95)
The right-hand side is not a tau function.
One can check that the graph in Fig 2c yields us the following matrix model
$\int\prod_{a=1}^{3}e^{\sum_{m>0}N\frac{p^{(a)}_{m}}{m}\mathrm{Tr}\left({\mathfrak{F}_{a}(Z)}\right)^{m}}d\mu(Z)=\sum_{\lambda\atop\ell(\lambda)\leq N}N^{-2|\lambda|}\frac{s_{\lambda}(\mathfrak{V})\prod_{a=1}^{3}s_{\lambda}(N\mathbf{p}^{(a)})}{s_{\lambda}(\mathbf{p}_{\infty})^{2}}$ (96)
where $\mathfrak{V}=C_{1}C_{-1}C_{2}C_{-2}$ and where $\mathfrak{F}_{1}(Z)=Z_{1}C_{1},\,\mathfrak{F}_{2}(Z)=Z_{2}C_{2},\,\mathfrak{F}_{2}(Z)=Z_{1}^{\dagger}C_{-1}Z_{2}^{\dagger}C_{-2}$. Under the constraints discussed in sections 2.2.1 and 2.2.2, the right-hand side (96) can be equated to the tau function of the KP and TL hierarchies.